OpenStudy (anonymous):

A cube has a side length of 7 feet. What is the volume of the cube? A. 21 ft3 B. 84 ft3 C. 294 ft3 D. 343 ft3 @Broskishelleh @dan815 @iambatman @tanya123

OpenStudy (anonymous):

\[V = a^3 \]

OpenStudy (anonymous):

So with the given formula your answer is...?

OpenStudy (broskishelleh):

It is D

OpenStudy (anonymous):

ohhh ok

OpenStudy (broskishelleh):

Do you have any more questions?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

A cube has a side length of 12 meters. What is the volume of the cube? A. 36 m3 B. 144 m3 C. 864 m3 D. 1728 m3

OpenStudy (anonymous):

Same thing, can you set it up for me this time, and @Broskishelleh please don't give the answer.

OpenStudy (anonymous):

ok nvm now i now how to do it

OpenStudy (anonymous):

it's D again

OpenStudy (broskishelleh):

D, just as I am Batman said before V = a^3 So V = 12 x 12 x 12